OpenStudy (anonymous):

College Algebra: reviewing for my final exam!! The growth in the population of a biology experiment fits A(t)=445e^0.013t where t is the number of years since 1987. Estimate the population in the year 2000. ** I tried doing it myself, but I'm not getting the correct answer**

OpenStudy (p0sitr0n):

you have to find t: t=2000-1987=13 thus A(13)=445 e^(0.013*13) =~ 527

OpenStudy (anonymous):

I got 13, but how did you solve it @P0sitr0n ? Step-by-step

OpenStudy (anonymous):

As far as I got on mine was \[A(13)=445e ^{0.169}\]

OpenStudy (p0sitr0n):

i guess using a calculator is mandatory for this lol

OpenStudy (anonymous):

I am only allowed to use a scientific calculator for the exam.

OpenStudy (p0sitr0n):

if you do not have e on your calculator, use e~ 2.718

OpenStudy (anonymous):

How about ln?

OpenStudy (p0sitr0n):

ln is the inverse function of e

OpenStudy (anonymous):

Okay, so what did you press on your calculator (in order) to get the answer?

OpenStudy (p0sitr0n):

\[e^x =y => \ln y =x\]

OpenStudy (p0sitr0n):

i have a button which is \[e^x\]

OpenStudy (anonymous):

Oh wait, I think I did it!

OpenStudy (anonymous):

I pressed 0.169, then e^x times 445 and got 526.6 which is 527

OpenStudy (p0sitr0n):

it may easier to memorize that \[e = \sum_{k=0}^{\infty} \frac{ 1 }{ k! } = \frac{ 1 }{ 1} + \frac{ 1 }{ 1 \times 2} + \frac{ 1 }{ 1 \times 2 \times 3} + ..\]

OpenStudy (p0sitr0n):

yeah thats the thing

OpenStudy (anonymous):

Thank you very much P0sitr0n!! I'll definitely remember this for my exam on Tuesday!! :)

OpenStudy (p0sitr0n):

just forget the big sum above and use the e^x button since you have it lol

OpenStudy (anonymous):

I will!! :)